OpenStudy (anonymous):

college question: when jack started his job working for an industrial manufacturing company, he contributed $211 at the end of each month into a savings account earned 2.4% interest compounded monthly for 7 years. at the end of the 7th year, jack was laid off. to help meet family expenses, jack withdrew $295 from the savings account at the end of each month for 2 years. At the end of the second year of being unemployed, jack found another job and started contributing $138 back into the savings account at the end of each month for the next six years. How much money would he have in the account

OpenStudy (anonymous):

I would solve this problem in three parts 1)First job 2)laid off 3) second job

OpenStudy (anonymous):

I will also assume that he started the account with 0 balance

OpenStudy (anonymous):

ok ill give it a try on the tvm solver

OpenStudy (radar):

Post your results

OpenStudy (anonymous):

im stuck. im trying to find out how to the laid off part. i first tried to just do 24 x 295 and that would be the amount that he took out total, but at the same time the account is earning interest so im not sure what to do

OpenStudy (anonymous):

wow this is really complicated lol

OpenStudy (anonymous):

Deposit= D R = Rate First Month - D Second Month - D+D(1+R) = D(1+(1+R)) Third Month - D(1+(1+R)+(1+R)^2) We see a pattern Nth Month - D(1+(1+R)+(1+R)^2+(1+R)^3+.....+(1+R)^N-1) You can add them up to 7 years but there is a short cut. (1+y+....+y^(n-1))(1-y)= (1-y^n) Using it on our equation.. D(1+(1+R)+(1+R)^2+(1+R)^3+.....+(1+R)^N)= D [ (1-(1+R)^N)/(1-(1-R))] D(-(1-(R+1)^N)/R)

OpenStudy (anonymous):

when i used the tvm solver i got n= 84 (12)(7) I%= 2.4 pv= 0 pmt= -211 and he would end up having 19278.88 before he got laid off?

OpenStudy (anonymous):

"he contributed $211 at the end of each month into a savings account earned 2.4% interest compounded monthly for 7 years." Let plug in number D(-(1-(R+1)^N)/R) 211(-(1-(.024/12+1)^63)/(.024/12))=$19278 That's how much he had before he got laid off

OpenStudy (anonymous):

so you were wrong lol thats not the part i needed help with i needed help with the part where here withdrew 295

OpenStudy (anonymous):

Here is formula, to avoid confusion I won't include the derivation(let me know if you're interested) P= Balance R= Interest( For Month) N= Number of Month W- Withdrawal \[P(1+R)^N-w\left(\frac{\left(1-(1+R)^N\right)}{-R}\right)\]

OpenStudy (anonymous):

Plugging in \[19278\left(1+\left(\frac{.024}{12}\right)\right)^{24}-295\left(\frac{\left(1-\left(1+\left(\frac{.024}{12}\right)\right)^{24}\right)}{-\left(\frac{.024}{12}\right)}\right)\] I got 12979.7

OpenStudy (anonymous):

Excel agrees with me

OpenStudy (anonymous):

Any question?

OpenStudy (anonymous):

yes if he withdrew 295 for 2 years, how is there more money in the account than before?

OpenStudy (anonymous):

Before there was $19278 after two years of withdrawal there is $12979.7; certainly no more than before.

OpenStudy (anonymous):

Any other?

OpenStudy (anonymous):

ohoh okay thanks i read it wrong ill go ahead and try to finish the last step now

OpenStudy (radar):

Thanks for posting, a real interesting problem (and solution)